OpenStudy (anonymous):

Find the length of the base of a square pyramid if the volume is 576 cubic inches and has a height of 3 inches. Use 3.14 for pi and round your answer to the nearest hundredth. a. 4 inches b. 8 inches c. 16 inches d. 24 inches

OpenStudy (e.mccormick):

Well, know the formula for the volume of a square pyramid?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got b for the answer but is it right?

OpenStudy (anonymous):

my teacher graded my homework and said it was wrong.

OpenStudy (e.mccormick):

So \(V=\dfrac{1}{3}b^2h\) You are given \(h=3\) and \(V = 576\) \(576=\dfrac{1}{3}b^23\)

OpenStudy (e.mccormick):

So when you solve that for b, what do you get?

OpenStudy (anonymous):

a?

OpenStudy (e.mccormick):

\(576=\dfrac{1}{3}b^23~ \leftarrow \) I mean that b. As in the length or height of the base, since on a square they are the same.

OpenStudy (anonymous):

oh so it is 24?

OpenStudy (e.mccormick):

Yah. The 1/3 and 3 cancel and you are left with finding the root of 576 or 24.